please help
find the value of y if 13.56y-14.76=3(4.12y-6.72)

Answer by checkley77(12844)

13.56y-14.76=3(4.12y-6.72)
13.56y-14.76=12.36y-20.16
13.56y-12.36y=14.76-20.16
1.2y=-5.4
y=-5.4/1.2
y=-4.5 answer.
proof:
13.56*-4.5-14.76=3(4.12*-4.5-6.72)
-61.02-14.76=3(-18.54-6.72)
-75.78=3*-25.26
-75.78=-75.78